OpenStudy (anonymous):

Solve the following equation. -4/13x=-20

OpenStudy (anonymous):

-4/13x=-20 take 13x over (multiply) -4 = -20(13x) -4 = -260x therefore: x = -4/-260 x = 1/65 :)